orthogonal trajectories of y' = - (x/y)


1
Expert's answer
2022-01-07T14:26:30-0500

"y' = -\dfrac{x}{y} \\
\text{To get the orthogonal trajectory}\\
\text{We replace} \, -\dfrac{x}{y} \, \text{by its negative reciprocal}\\
\text{Hence, we have}\\
y' = \dfrac{y}{x} \\
\text{We solve this D.E using separation of variables}\\
\text{Hence, we have}\\
\dfrac{dy}{dx} = \dfrac{y}{x} \\
\text{Integrating both sides, we have}\\
\int \dfrac{dy}{y} = \int \dfrac{dx}{x} \\
\text{Integrating, we obtain}\\
y = cx \\
\text{which is a family of the orthogonal trajectory}\\
\text{where C is an arbitrary constant}"